There is a 9.2 g piece of gold. What volume will the gold have? (Density of gold = 19.4 g/cm^3) round answer to tenths place (1 decimal place)
1
Expert's answer
2020-10-23T09:16:28-0400

V=m/p

V=9,2/19,4=0,474=0,5

Answer 0,5 см3